Problem solving: The problem of two positive numbers is 245. Determine the two number if one number is 5 times the other.
Answer by mananth(16946)

let one number bex
the other number = 5x
product = 5x^2
5x^2= 245
x^2= 245/5
x^2=49
x=7
other number = 35